OpenStudy (anonymous):

Solve. Write the answer in simplest form. (Must show work!) 1. 3/4m = 5/6

OpenStudy (arbershabani97):

3/4m=5/6 /*4 3m=20/6 / /3 (20/6=10/3) m=10/3 / 3 when you can't divide you change the sign to multiply you change the number from 3/1 to 1/3 so m=10/3 * 1/3 m=10/9

OpenStudy (anonymous):

Isnt the answer is m = 1 1/9 ? @arbershabani97

OpenStudy (arbershabani97):

its the same 1 1/9 with 10/9

OpenStudy (anonymous):

@arbershabani97 Does the "/" sign means to divide or just fractions ? I'm confused.

OpenStudy (arbershabani97):



OpenStudy (anonymous):

So fractions ?

OpenStudy (anonymous):

of all of them?

OpenStudy (mathlegend):

@Mertsj this is what I mean...

OpenStudy (mathlegend):

Since this question is closed and is a repeat. I just want to ask @connor_ do you really understand?

OpenStudy (anonymous):

Yes @MathLegend